# The AIDA-TNG project: dark matter profiles and concentrations in alternative dark matter models

### v1

@@
In the standard Cold Dark Matter (CDM) scenario, the density profiles of dark matter haloes are well described by analytical models linking their concentration to halo mass. Alternative scenarios, such as warm dark matter (WDM) and self-interacting dark matter (SIDM), modify the inner structure of haloes and predict different profile shapes and central slopes. We employ the AIDA-TNG simulations to investigate how alternative dark matter physics and baryonic processes jointly shape the internal structure of haloes. Using dark-matter-only and full-physics runs, we measure the dark matter density profiles of haloes spanning six orders of magnitude in mass, from $10^{9.5}M_{\odot}$ to $10^{14.5}M_{\odot}$, and characterise them with multiple analytical models. We provide the distribution of the best-fitting parameters, as well as the concentration-mass relation in WDM and SIDM. The Einasto profile well reproduces the inner flattening produced in WDM models, both in the collisionless and in the full-physics runs. In SIDM dark-matter-only runs, haloes are better described by explicitly cored profiles, with core sizes that depend on mass and on the self-interaction model. When baryons are included, the differences between CDM and SIDM decrease, and such large dark-matter cores no longer form because adiabatic contraction in the baryon-dominated region counteracts self-interactions. Nevertheless, the coupling between baryons and self-interactions induces a broader range of inner slopes, including cases that are steeper than CDM at Milky Way masses. Alternative dark matter physics thus leaves clear signatures in the inner halo structure, even if baryons significantly reshape these differences. Our results are useful for future studies that need to predict the properties of haloes in multiple dark matter models.
@@ 1 Introduction
The Cold Dark Matter (CDM) model assumes that dark matter consists of heavy, collisionless particles interacting only through gravity. In this framework, the formation and evolution of cosmic structures are well described by hierarchical clustering, and the internal structure of haloes arises naturally from gravitational collapse (White & Rees, 1978; Blumenthal et al., 1984; Davis et al., 1985). In particular, the spherically averaged density profiles of dark matter haloes are well represented by the classic Navarro–Frenk–White (NFW) model (Navarro et al., 1996, 1997) or by the Einasto profile (Merritt et al., 2006; Navarro et al., 2004; Springel et al., 2008). The compactness of the density distribution is commonly quantified through the concentration-mass relation. This relation has been extensively calibrated in collisionless simulations (Dolag et al., 2004; Duffy et al., 2008; Giocoli et al., 2012b; Ludlow et al., 2013, 2016; Dutton & Macciò, 2014), showing that more massive haloes are less concentrated than less massive ones and that concentrations decrease with redshift (Bullock et al., 2001; Wechsler et al., 2002). Hydrodynamical simulations have demonstrated that the introduction of baryonic processes significantly modifies this simple picture. The current generation of high-resolution hydrodynamical simulations (Schaye et al., 2015; Pillepich et al., 2018a; Nelson et al., 2018; Davé et al., 2019) now models galaxy formation with sufficient realism to reproduce key properties of galaxies and their satellites over many orders of magnitude in mass. The processes involved, such as gas cooling, star formation, feedback, and dynamical interactions, can either steepen the central density profile through adiabatic contraction (Blumenthal et al., 1986; Gnedin et al., 2004) or produce central cores through energetic feedback (Governato et al., 2010; Pontzen & Governato, 2012). As a result, the linear dependence between concentration and mass can be altered or even broken (Diemer & Kravtsov, 2015; Sorini et al., 2025).
@@
The introduction of baryonic physics in simulations, together with a parallel improvement in the observational domain, has also contributed to solving some long-standing tensions between CDM predictions and observations, such as the missing satellite problem (Klypin et al., 1999; Drlica-Wagner et al., 2015; Kim et al., 2018; Wetzel et al., 2016) or the formation of cored density profiles via feedback processes. However, other small-scale problems are still unsolved in CDM, as the diversity of galaxy rotation curves (Oman et al., 2015; Creasey et al., 2017) and the abundance of over-concentrated systems detected via gravitational lensing (Meneghetti et al., 2020, 2023; Minor et al., 2021; Despali et al., 2025a; Enzi et al., 2024). Alternative dark matter models, such as warm (WDM) and self-interacting (SIDM) dark matter, are candidates to explain these discrepancies since they tend to modify the number density and density structure of haloes. The primary effect of WDM is to suppress the number of low-mass haloes compared to CDM, but the high velocity of particles naturally also lowers their concentration, producing a turnover in the concentration-mass relation (Lovell et al., 2014; Ludlow et al., 2016; Bose et al., 2016). In SIDM, non-gravitational interactions between particles lead to the exchange of energy and momentum in high-density regions and thus to the creation of a central density core, flatter than that seen in WDM and present at all halo masses (Vogelsberger et al., 2012, 2016). The size of this central core depends on the frequency of the interactions between dark matter particles, regulated by the self-interaction cross section: in dark-matter-only (DMO) simulations, a larger cross section corresponds to larger density cores. However, when the cross section is very high ($\sigma/m_{\chi}\geq$ 100 cm2/g), a fast core formation can be followed by a second phase of gravothermal collapse towards the centre, leading to a rapid increase of the central density (Tran et al., 2025; Shah & Adhikari, 2024; Yang et al., 2023; Kong et al., 2025).
@@
While most simulation projects are interested in one specific mass range, here we are interested in testing alternative dark matter models over a wide range of halo and galaxy types and environments. This is essential, not only to derive constraints and exclusion regions for alternative models, but also to provide scaling relations for WDM and SIDM that are calibrated as well as their CDM counterparts. In this way, we can ask if any other model can explain the process of structure formation and match astrophysical observations as well as CDM. The AIDA-TNG simulations (Despali et al., 2025b) are ideal for this purpose since they provide a consistent set of high-resolution cosmological boxes simulated with the same galaxy formation model and multiple WDM and SIDM models, in addition to CDM. In this work, we characterise the dark matter profiles of CDM, SIDM, and WDM haloes spanning six orders of magnitude in mass, from $10^{9.5}$ to $10^{14.5}M_{\odot}$. We compare the DMO and full-physics (FP) runs to study the interplay between baryons and dark matter variations, breaking the degeneracy between the two at small scales.
The paper is structured as follows. We present the AIDA-TNG runs and the models for the density profiles in Sect. 2 and Sect. 3, respectively. In Sect. 4, we start by discussing the impact of alternative models on the profiles in the DMO runs and then move to the combined effect of these with baryonic physics in Sect. 5, addressing the difference in adiabatic contraction between CDM, WDM, and SIDM. In both cases, we model the dark matter density distribution with multiple analytical models and derive the concentration-mass relation at $z=0$. In Sect. 6, we discuss how the profile properties and best-fit parameters vary with redshift. Finally, we summarise the results and present our conclusions in Sect. 7. In the Appendices, we discuss more technical aspects related to the goodness-of-fit of different models (Appendix A) and resolution effects (Appendix. B).
@@ 2 The AIDA simulations
Table 1: Summary of the AIDA-TNG runs used in this work. [^1] We list the mass and spatial resolution of each run and mark the dark matter models that are used in this work. Those marked in grey are not used in the main body of the paper and only included for resolution convergence tests (see Appendix B).
@@
Haloes and subhaloes are identified using the SUBFIND algorithm, producing catalogues of the halo and galaxy properties with the same information available for the IllustrisTNG runs (Nelson et al., 2019). Throughout the paper, we define halo masses as $M_{200c}$, corresponding to the mass inside the radius $R_{200c}$ where the enclosed density equals 200 times the critical density at the redshift of the considered snapshot. Masses and lengths are given in units of $M_{\odot}$ and kpc, respectively.
@@ 3.2 Analytical models
We fit the dark matter density profiles with all analytical models, independently of the underlying dark matter physics, in order to evaluate which parametrisation best describes CDM, WDM, and SIDM haloes. To ensure the robustness of the fits, we restrict the analysis to radial bins in the range ($3\epsilon_{\rm DM}\leq r\leq r_{200c}$) (see Fig. 3), where $\epsilon_{\rm DM}$ is the Plummer-equivalent softening length; the lower limit excludes the innermost region affected by numerical two-body relaxation and force softening. This means that we consider $M_{\rm 200c}\geq 5\times 10^{9}M_{\odot}$ in the 50/A boxes and $M_{\rm 200c}\geq 7\times 10^{10}M_{\odot}$ in the 100/A and 50/B ones.
@@ 5.1 Adiabatic contraction in SIDM
Figure 6: Left: dark matter density profiles of haloes in the mass range $\log M_{200c}=[10^{12}-10^{13}]$ $M_{\odot}$ in the 50/A run, for the CDM (left), vSIDM (middle) and SIDM1 (right) DMO (top) and FP (bottom) runs. We only show the profiles for $r\geq\epsilon_{DM}$ and the vertical dashed line marks $r=3\epsilon_{DM}$, while the grey band shows is the range used to measure $\gamma_{DM}$. Right: Inner slope of the dark matter density profiles from the left panel, plotted as a function of halo mass and coloured by the stellar half-mass radius r∗,1/2 of the central galaxy.
@@ 5.2 Density profiles and concentrations
The steepening of the inner profile suppresses the development of SIDM-induced constant-density cores in the FP runs. As a result, the Y24 core radius becomes very small at all masses, and the Y24 functional form loses its advantage over Einasto: as confirmed by the $\chi^{2}$ distributions reported in Appendix A, the Einasto model becomes the preferred description even for SIDM1 and vSIDM haloes, except at the low-mass end $_{200c}\leq 10^{11.5}MM_{\odot}$. As discussed in the previous section, the FP SIDM1 haloes in the mass range $M_{200}=[10^{12},\,10^{13}]{M_{\odot}}$ show a particularly wide diversity of inner slopes, including systems with very steep cusps ($\gamma<-1.5$). This is evident by the peak in scale density and the corresponding dip in scale radius for both profiles. The same scale also shows an increase in concentration, as we see in the bottom panels of Fig. 7. Here we show the concentration–mass relation for the FP CDM run and compare it with previous TNG-based studies. We reproduce the results of Sorini et al. (2025), who found that TNG baryonic physics produces a sharp enhancement in concentration around $M_{200c}\simeq 10^{12}M_{\odot}$. Remarkably, the same feature appears in all dark matter models, demonstrating that the baryonic impact on halo structure is largely independent of the underlying dark matter microphysics at this mass scale. This mass range also corresponds to the maximum of the stellar-to-halo mass relation of the AIDA runs (Despali et al., 2025b), in agreement with earlier findings that variations in concentration correlate tightly with the stellar mass fraction (Di Cintio et al., 2014a, b). In Fig. 8, we show this connection explicitly: the peak of the stellar mass fraction mirrors the peak in the ratio of FP-to-DMO mean concentrations, confirming a direct link between baryonic assembly efficiency and the strength of adiabatic contraction effects.
@@ 7 Summary and conclusions
In this work, we investigate the dark matter density profile of haloes in the AIDA-TNG simulations (Despali et al., 2025b), considering systems between $5\times 10^{9}M_{\odot}$ and $5\times 10^{14}M_{\odot}$. This analysis is the first to address models for the density profiles in CDM, WDM, and SIDM models over such a wide range of halo masses, including the effects of full physical baryon processes. This is possible thanks to the setup of the AIDA-TNG sample that includes cosmological hydrodynamical simulations of the same boxes, considering multiple alternative dark matter models: warm dark matter (WDM) with particle masses of 1 and 3 keV (WDM1, WDM3), and self-interacting dark matter with either a constant (SIDM1) or velocity-dependent (vSIDM) scattering cross-section. We measure the halo dark matter profiles in all runs, study the best-fit analytical models for each scenario, and examine the concentration-mass relation with and without the inclusion of baryons.
In Sect. 4, we first look at the DMO runs, in a way to isolate the impact of the dark matter physics on halo structures. In WDM models, the dark matter profiles are well described by the Einasto model (Einasto, 1965), with no particular evidence of central cores. The suppression of small-scale power leads to systematically lower concentrations at low masses consistent with the results of previous works (Lovell et al., 2014; Ludlow et al., 2016). On the other side, the DMO SIDM models, with constant (SIDM1) and velocity-dependent (vSIDM) cross-section: the dark matter profiles in the DMO runs form large and flat inner cores, which are best described by an analytical model which explicitly includes a core, such as the Y24 profile (Yang et al., 2024). The core size increases with halo mass but presents a different dependence on mass in the two models: vSIDM haloes form relatively larger cores in low-mass systems (due to higher cross-section at low velocities), while SIDM1 haloes produce larger cores at the high-mass end. It is important to underline that in both cases, the size of the core radius remains smaller than the scale radius of the profile, meaning that the core does not affect halo concentrations, which are essentially unchanged from CDM as shown in Fig. 4. This implies that a complete description of the structure of SIDM haloes requires combining the concentration-mass relation with a model for the core size (see Eq. 3).

### v2

@@
In the standard Cold Dark Matter (CDM) scenario, the density profiles of dark matter haloes are well described by analytical models linking their concentration to halo mass. Alternative scenarios, such as warm dark matter (WDM) and self-interacting dark matter (SIDM), modify the inner structure of haloes and predict different profile shapes and central slopes. We employ the AIDA-TNG simulations to investigate how alternative dark matter physics and baryonic processes jointly shape the internal structure of haloes. Using dark-matter-only and full-physics runs, we measure the dark matter density profiles of haloes spanning six orders of magnitude in mass, from $10^{9.5}M_{\odot}$ to $10^{14.5}M_{\odot}$, considering radial bins that are well-resolved above the spatial resolution of the simulations ($r\geq 4.4$ kpc and $r\geq 1.7$ kpc for the 110.7 and 51.7 Mpc boxes, respectively). We fit the profiles with multiple analytical models and provide the distribution of the best-fitting parameters, as well as the concentration-mass relation in WDM and SIDM. The Einasto profile well reproduces the inner flattening produced in WDM models, both in the collisionless and in the full-physics runs. In SIDM dark-matter-only runs, haloes are better described by explicitly cored profiles, with core sizes that depend on mass and on the self-interaction model. When baryons are included, the differences between CDM and SIDM decrease, and such large dark-matter cores no longer form because adiabatic contraction in the baryon-dominated region counteracts self-interactions. Nevertheless, the coupling between baryons and self-interactions induces a broader range of inner slopes, including cases that are steeper than CDM at Milky Way masses. Alternative dark matter physics thus leaves clear signatures in the inner halo structure, even if baryons significantly reshape these differences. Our results are useful for future studies that need to predict the properties of haloes in multiple dark matter models.
@@ 1 Introduction
The Cold Dark Matter (CDM) model assumes that dark matter consists of heavy, collisionless particles interacting only through gravity. In this framework, the formation and evolution of cosmic structures are well described by hierarchical clustering, and the internal structure of haloes arises naturally from gravitational collapse (White & Rees, 1978; Blumenthal et al., 1984; Davis et al., 1985). In particular, the spherically averaged density profiles of dark matter haloes are well represented by the classic Navarro–Frenk–White (NFW) model (Navarro et al., 1996, 1997) or by the Einasto profile (Merritt et al., 2006; Navarro et al., 2004; Springel et al., 2008). The compactness of the density distribution is commonly quantified through the concentration-mass relation. This relation has been extensively calibrated in collisionless simulations (Dolag et al., 2004; Duffy et al., 2008; Giocoli et al., 2012b; Ludlow et al., 2013, 2016; Dutton & Macciò, 2014), showing that more massive haloes are less concentrated than less massive ones and that concentrations decrease with redshift (Bullock et al., 2001; Wechsler et al., 2002). Hydrodynamical simulations have demonstrated that the introduction of baryonic processes significantly modifies this simple picture. The current generation of high-resolution hydrodynamical simulations (Schaye et al., 2015; Pillepich et al., 2018a; Nelson et al., 2018; Davé et al., 2019) now models galaxy formation with sufficient realism to reproduce key properties of galaxies and their satellites over many orders of magnitude in mass. The processes involved, such as gas cooling, star formation, feedback, and dynamical interactions, can either steepen the central density profile through adiabatic contraction (Blumenthal et al., 1986; Gnedin et al., 2004) or produce central cores through energetic feedback (Governato et al., 2010; Pontzen & Governato, 2012). As a result, the linear dependence between concentration and mass can be altered or even broken (Diemer & Kravtsov, 2015; lovell18; anbajagane22; Sorini et al., 2025).
@@
The introduction of baryonic physics in simulations, together with a parallel improvement in the observational domain, has also contributed to solving some long-standing tensions between CDM predictions and observations, such as the missing satellite problem (Klypin et al., 1999; Drlica-Wagner et al., 2015; Kim et al., 2018; engler21; Wetzel et al., 2016) or the formation of cored density profiles via feedback processes. However, other small-scale problems are still unsolved in CDM, as the diversity of galaxy rotation curves (Oman et al., 2015; Creasey et al., 2017) and the abundance of over-concentrated systems detected via gravitational lensing (Meneghetti et al., 2020, 2023; Minor et al., 2021; Despali et al., 2025a; Enzi et al., 2024). Alternative dark matter models, such as warm (WDM) and self-interacting (SIDM) dark matter, are candidates to explain these discrepancies since they tend to modify the number density and density structure of haloes. The primary effect of WDM is to suppress the number of low-mass haloes compared to CDM, but the high velocity of particles naturally also lowers their concentration, producing a turnover in the concentration-mass relation (Lovell et al., 2014; Ludlow et al., 2016; Bose et al., 2016). In SIDM, non-gravitational interactions between particles lead to the exchange of energy and momentum in high-density regions and thus to the creation of a central density core, flatter than that seen in WDM and present at all halo masses (Vogelsberger et al., 2012, 2016). The size of this central core depends on the frequency of the interactions between dark matter particles, regulated by the self-interaction cross section: in dark-matter-only (DMO) simulations, a larger cross section corresponds to larger density cores. However, when the cross section is very high ($\sigma/m_{\chi}\geq$ 100 cm2/g), a fast core formation can be followed by a second phase of gravothermal collapse towards the centre, leading to a rapid increase of the central density (Tran et al., 2025; Shah & Adhikari, 2024; Yang et al., 2023; Kong et al., 2025).
@@
While most simulation projects are interested in one specific mass range, here we aim at testing alternative dark matter models over a wide range of halo and galaxy types and environments. This is essential, not only to derive constraints and exclusion regions for alternative models, but also to provide scaling relations for WDM and SIDM that are as well calibrated as their CDM counterparts. In this way, we can ask if any other model can explain the process of structure formation and match astrophysical observations as well as CDM. The AIDA-TNG simulations (Despali et al., 2025b) are ideal for this purpose since they provide a consistent set of high-resolution cosmological boxes simulated with the same galaxy formation model and multiple WDM and SIDM models, in addition to CDM. In this work, we characterise the dark matter profiles of CDM, SIDM, and WDM haloes spanning six orders of magnitude in mass, from $10^{9.5}$ to $10^{14.5}M_{\odot}$. We compare the DMO and full-physics (FP) runs to study the interplay between baryons and dark matter variations, breaking the degeneracy between the two at small scales.
The paper is structured as follows. We present the AIDA-TNG runs and the models for the density profiles in Sect. 2 and Sect. 3, respectively. In Sect. 4, we start by discussing the impact of alternative models on the profiles in the DMO runs and then move to the combined effect of these with baryonic physics in Sect. 5, addressing the difference in adiabatic contraction between CDM, WDM, and SIDM. In both cases, we model the dark matter density distribution with multiple analytical models and derive the concentration-mass relation at $z=0$. In Sect. 6, we discuss how the profile properties and best-fit parameters vary with redshift. Finally, we summarise the results and present our conclusions in Sect. 7. In the Appendices, we discuss more technical aspects related to the goodness-of-fit of different models (Appendix A) and resolution effects (Appendix B).
@@ 2 The AIDA simulations
Table 1: Summary of the AIDA-TNG runs used in this work. [^1] We list the mass and spatial resolution of each run and mark the dark matter models that are used in this work. We use $\epsilon_{DM,*}$, the Plummer-equivalent softening length, to quantify the spatial resolution of each run. Those marked in grey are not used in the main body of the paper and only included for resolution convergence tests (see Appendix B).
@@
Haloes and subhaloes are identified using the SUBFIND algorithm, producing catalogues of the halo and galaxy properties with the same information available for the IllustrisTNG runs (Nelson et al., 2019). Throughout the paper, we define halo masses as $M_{200c}$, corresponding to the mass inside the radius $R_{200c}$ where the enclosed density equals 200 times the critical density at the redshift of the considered snapshot. Masses and lengths are given in units of $M_{\odot}$ and comoving kpc, respectively.
@@ 3.2 Analytical models
We fit the dark matter density profiles with all analytical models, independently of the underlying dark matter physics, in order to evaluate which parametrisation best describes CDM, WDM, and SIDM haloes. To ensure the robustness of the fits, we restrict the analysis to radial bins in the range ($3\epsilon_{\rm DM}\leq r\leq r_{200c}$) (see Fig. 3), where $\epsilon_{\rm DM}$ is the Plummer-equivalent softening length; the lower limit excludes the innermost region affected by numerical two-body relaxation and force softening. This means that we consider $M_{\rm 200c}\geq 5\times 10^{9}M_{\odot}$ and $r\geq 1.7$ kpc in the 50/A boxes, while $M_{\rm 200c}\geq 7\times 10^{10}M_{\odot}$ and $r\geq 4.4$ kpc in the 100/A and 50/B ones.
@@ 5.1 Adiabatic contraction in SIDM
Figure 6: Left: dark matter density profiles of haloes in the mass range $\log M_{200c}=[10^{12}-10^{13}]$ $M_{\odot}$ in the 50/A run, for the CDM (black), vSIDM (green) and SIDM1 (blue) FP runs. We only show the profiles for $r\geq\epsilon_{DM}$ and the vertical dashed line marks $r=3\epsilon_{DM}$, while the grey band shows is the range used to measure $\gamma_{DM}$. Right: Inner slope of the dark matter density profiles from the left panel, plotted as a function of halo mass for the three runs and coloured by the stellar half-mass radius r∗,1/2 of the central galaxy.
@@ 5.2 Density profiles and concentrations
The steepening of the inner profile suppresses the development of SIDM-induced constant-density cores in the FP runs. As a result, the Y24 core radius becomes very small at all masses, and the Y24 functional form loses its advantage over Einasto: as confirmed by the $\chi^{2}$ distributions reported in Appendix A, the Einasto model becomes the preferred description even for SIDM1 and vSIDM haloes, except at the low-mass end $_{200c}\leq 10^{11.5}MM_{\odot}$. As discussed in the previous section, the FP SIDM1 haloes in the mass range $M_{200}=[10^{12},\,10^{13}]{M_{\odot}}$ show a particularly wide diversity of inner slopes, including systems with very steep cusps ($\gamma<-1.5$). This is evident by the peak in scale density and the corresponding dip in scale radius for both profiles. The same scale also shows an increase in concentration, as we see in the bottom panels of Fig. 7. Here we show the concentration–mass relation for the FP CDM run and compare it with previous TNG-based studies. We reproduce the results of previous works lovell18; anbajagane22; Sorini et al. (2025), who found that TNG baryonic physics produces a sharp enhancement in concentration around $M_{200c}\simeq 10^{12}M_{\odot}$. Remarkably, the same feature appears in all dark matter models, demonstrating that the baryonic impact on halo structure is largely independent of the underlying dark matter microphysics at this mass scale. This mass range also corresponds to the maximum of the stellar-to-halo mass relation of the AIDA runs (Despali et al., 2025b), in agreement with earlier findings that variations in concentration correlate tightly with the stellar mass fraction (Di Cintio et al., 2014a, b). In Fig. 8, we show this connection explicitly: the peak of the stellar mass fraction mirrors the peak in the ratio of FP-to-DMO mean concentrations, confirming a direct link between baryonic assembly efficiency and the strength of adiabatic contraction effects.
@@ 7 Summary and conclusions
In this work, we investigate the dark matter density profile of haloes in the AIDA-TNG simulations (Despali et al., 2025b), considering systems with total halo mass between $5\times 10^{9}M_{\odot}$ and $5\times 10^{14}M_{\odot}$. This analysis is the first to address models for the density profiles in CDM, WDM, and SIDM models over such a wide range of halo masses, including the effects of full physical baryon processes. This is possible thanks to the setup of the AIDA-TNG sample that includes cosmological magneto-hydrodynamical simulations of the same boxes, considering multiple alternative dark matter models: warm dark matter (WDM) with particle masses of 1 and 3 keV (WDM1, WDM3), and self-interacting dark matter with either a constant (SIDM1) or velocity-dependent (vSIDM) scattering cross-section. We measure the halo dark matter profiles in all runs, study the best-fit analytical models for each scenario, and examine the concentration-mass relation with and without the inclusion of baryons. To ensure the robustness of the results, we only trust radial bins $r\geq 3\epsilon_{\rm DM}$, where $\epsilon_{\rm DM}$ is the Plummer-equivalent softening length that is used to quantify the simulation resolution. This corresponds to $r\geq 4.4$ kpc for the 100/A runs and $r\geq 1.7$ kpc for the 50/A runs.
In Sect. 4, we first look at the DMO runs, in a way to isolate the impact of the dark matter physics on halo structures. In WDM models, the dark matter profiles are well described by the Einasto model (Einasto, 1965), with no particular evidence of central cores. The suppression of small-scale power leads to systematically lower concentrations at low masses consistent with the results of previous works (Lovell et al., 2014; Ludlow et al., 2016). On the other side, in the DMO SIDM models, with constant (SIDM1) and velocity-dependent (vSIDM) cross-section, the dark matter profiles form large and flat inner cores, which are best described by an analytical model which explicitly includes a core, such as the Y24 profile (Yang et al., 2024). The core size increases with halo mass but presents a different dependence on mass in the two models: vSIDM haloes form relatively larger cores in low-mass systems (due to higher cross-section at low velocities), while SIDM1 haloes produce larger cores at the high-mass end. It is important to underline that in both cases, the size of the core radius remains smaller than the scale radius of the profile, meaning that the core does not affect halo concentrations, which are essentially unchanged from CDM as shown in Fig. 4. This implies that a complete description of the structure of SIDM haloes requires combining the concentration-mass relation with a model for the core size (see Eq. 3).